Subject: APO 3.5m users committee minutes, February 2014

From: Suzanne Hawley

Submitted: Sun, 16 Feb 2014 11:27:55 -0800 (PST)

Message number: 1326 (previous: 1325, next: 1327 up: Index)

 		APO 3.5-m Users Committee Phonecon, 2/10/14

Attending: Klaene, Ketzeback, Gillespie, Bally, Kuzio de Naray, Kaiser, 
Holtzman, Strauss, Indebetouw, Anderson, Hawley

Absent: McMillan, Huehnerhoff, Bary, Walterbos, Hughes Clark, Harper, 


User feedback and comments from institutional representatives:

o Washington - nothing to report.

o Chicago - no report.

o New Mexico State - everything going fine.

o Princeton - everyone is happy.

o Colorado - nothing to report.

o Virginia - nothing to report.

o Johns Hopkins - Mary Beth thanked Suzanne for her recent visit to JHU. 
Suzanne had a town hall meeting with Hopkins users, which was well 

o Georgia State - Rachel reported that users at GS have not experienced 
problems reducing data with new timestamp formats in the headers. She also 
reported that her first night of remote observations went well.

o Seattle University - no report.

o NAPG - no report.

o University of Oklahoma - no report.


Discussion of telescope/instruments report:

In addition to the normal telescope report Mark mentioned a problem with 
one of the M2 actuators. It would usually happen at the first move of the 
night after defining the instrument change. The problem could not be 
reproduced by the engineering crew during the day time. The observing 
specialists have developed a quick recovery method to minimize downtime if 
this occurs again. The problem has not reproduced itself in a couple of 

Mark Klaene's detailed report for the reporting period follows:


 	      3.5-m Telescope and Instruments Highlights
 	                01/08/14 through 02/05/14

0) Overview

Generally a productive period.  Unseasonably high winds hampered observing 
at times.

1) Telescope

Telescope has been working very well.  There has been good progress made 
on the guider rewrite and some observers have volunteered to test the new 
guiding software on sky since engineering time is limited this time of 
year.  We will continue to make progress on this over the next several 
months and when satisfied convert the new guider routines to be the 

2) Instruments

DIS has been operating normally after the grating bay limit switch was 
repaired.  We also characterized and added the B1000 grating to our 
inventory.  DCAM camera replacement has been delayed (Apogee Inc. was 
bought out and that has delayed their delivery).  Commissioning will be 
moved to the summer.

Agile is operational with no problems.

TripleSpec had no operational problems this period.

SPIcam variable dark current has been isolated to the current ion pump 
type and controller.  We continue to work with the ION pump manufacturer 
to come up with a long term solution. bCurrently we see a spike that may 
affect dark current about 1/day and if it occurs during observations we 
will let the observer know to check their data.  Generally the effect is 

Echelle is operational with no problems.

NIC-FPS is operating in shared-risk observing mode due to occasional image 
corruption issue.  NFS less than 9 is operational.


New users info, training:

A draft of a new users webpage was sent to the committee members before 
the meeting.  Suzanne described the motivation as a road map for new users 
rather than a reproduction of all the material already found on the site. 
Michael said it looks great but said some of the acronyms can be 
mysterious to new users. He made references to some brief examples 
(ARC,TUI, TCC) and asked that those be explained within the text of the 
page. Jon Bally wanted to know if it was going to be linked on the front 
APO page. Michael and Mary Beth seconded this as being important for new 
users to be able to find this page. Rachel suggested adding driving 
directions (or links) to the new users page. She said that the folks at 
GSU were able to locate much of the information on their own however some 
of the information is less straightforward to find such as directions to 
the site. Mention about not having a cafeteria on site and purchasing food 
and explanation of reservations to lodging should be added as well to the 
page. Remy mentioned that the term "non-standard observing" will not make 
sense to a new user and should be changed.

Suzanne thanked Alaina Bradley for putting the page together. She will 
work with Russet and Alaina to get these suggested changes made to the new 
users page, and we will then advertise it to the full users community.

APO users wiki:

Jon Holtzman has started an APO 3.5-m wiki page 
( for one and all to post on, and refer to, 
regarding most any operational and data analysis topic related to the 
3.5-m telescope and instruments.  The idea is for users to help each other 
by sharing experiences and lessons-learned.  Right now, Jon is undecided 
about posting and registration policies and procedures, e.g., we might 
make it read-only for unregistered users, but you would need to register 
to have writing privileges.  Suzanne added that everyone should be able to 
and encouraged to contribute to the wiki.  Jon encouraged the site staff 
to also add material to the wiki.

Bill said that the data reduction section is incomplete in many of the 
instrument manuals and suggested the wiki would be a good place for this 
information and could be linked from the manuals.  The new users page 
should also mention the wiki.  He will speak with Alaina Bradley about 
these changes.

Jon said that he would post a general notice about the wiki separately 
from the Users Committee minutes.  All users are encouraged to look at the 
wiki and to contribute if they have useful information.


Guider image matching scripts:

Bill recalled that we had discussed quite a while ago the desire to match 
guider frames with their corresponding science data exposures.  At that 
time, Bill modified a script originally developed by Adam Kowalski (UW) so 
that it was generalized to all instruments and guiders at the telescope. 
This feature is currently available on Newton, and is described on the 
user web pages.  But some of the scripts are now broken because of changes 
recently made to the FITS headers.  The issue is should we fix the scripts 
(which will take a few days of effort), and if so, at what priority. 
There followed a discussion that led to the following action:

===> ACTION:  If users feel that the guider matching scripts are 
important, let Bill and Mark know.  Also, if there is someone willing to 
volunteer to help fix the scripts, likewise let Bill and Mark know.

If no feedback is received, Mark and Bill said that they plan to fix the 
scripts over the coming months as time allows.


ARCSAT update:

Bill reported that the controller that enables remote dome operation is 
nearing completion, and will be tested in the coming days.  Several users 
will be involved with beta-testing remote operation in the next few weeks. 
Suzanne said that we plan to make ARCSAT available on a shared-risk, 
no-cost, basis during Q2.  John Bally asked if ARCSAT can be fitted with 
narrow-band filters, and Bill said he thought that the APO 2 x 2-inch 
filters will work; Suzanne thought there might be a thickness constraint 
that should be checked.

===> ACTION: Send expressions of interest in using ARCSAT during Q2 to 


GIFS training plan in Q2:

   Suzanne and Mark have had some discussions with Carol Grady and Michael 
McElwain (GSFC) regarding the availability and user training on the 
Goddard Integral Field Spectragraph (GIFS). Carol and Michael would like 
to hold an open GIFS training session for interested users in May/June 
during a dark run. GIFS will by then be fully integrated in TUI. Users 
interested in learning how to take and reduce data with the instrument are 
encouraged to propose 1 to 2 half nights for Q2 through their normal TAC 
during this time frame. Basic specifications and capabilities on the 
instrument are available on the APO instrument page but more details can 
be obtained by contacting Carol Grady. The successful proposals will then 
be block scheduled together so that users can travel to APO and learn to 
take and reduce GIFS data with the instrument team.


WIYN time in NOAO 2014B semester:

Suzanne said that we are experimenting with a small time trade (3 nights) 
with WIYN during the 2014A semester.  For the 2014B semester, Suzanne asks 
that if people are interested in trading APO 3.5-m nights with WIYN, one 
for one, they should send her an expression of interest so that she can 
work out the scheduling arrangements on an ad hoc basis.  Probably four or 
five full nights in total will be available.  The APO time would come out 
of Q3 or Q4 institutional allocations.  It was suggested that we might 
explore time trades with SOAR, and Suzanne agreed to look into this.

ACTION: Send expressions of interest for WIYN time trades to Suzanne


20th anniversary celebration:

This year marks the 20th anniversary of the dedication of the 3.5-m 
telescope, and the 30th anniversary of the founding of ARC.  Suzanne said 
that we have decided not to hold a large science meeting in celebration, 
but rather to hold an event on May 12 and 13 where some 50 people will 
meet at the site, view through the telescope, listen to talks largely on 
topics related to ARC and the 3.5-m telescope's history and usage, and 
celebrate later with dinner at the Lodge.  This is also an opportunity for 
new institutional users and their administrators to visit the site and 
learn about ARC.  We need suggestions from the users committee on who to 

ACTION: Send Suzanne your invitation list (if you have not done so yet) 



[open from previous meeting]:

 	===> ACTION: Committee members please query their users about 
issues with reducing data due to the new format of the timestamp.

              Status:  Discussed above, closed.

 	===> ACTION:  Users Committee members to think about the context 
and themes for a 20th-anniversary celebration for the 3.5-m, and discuss 
at the next Users Committee telecon.

              Status:  Discussed above, will leave one week more for 
comments, and invitation list needed ASAP.

[new from this meeting]:

 	===> ACTION:  If users feel that the guider matching scripts are 
important, let Bill and Mark know.  Also, if there is someone willing to 
volunteer to help fix the scripts, likewise let Bill and Mark know.

 	===> ACTION:  Send expressions of interest in using ARCSAT in Q2 
to Suzanne

         ===> ACTION:  Send expressions of interest for WIYN time trades to 


Next meeting:  The next Users Committee phonecon will be on Monday 10 
March at 8:30 AM Pacific Time.  The agenda and other materials will be 
sent to the committee members during the preceding week.

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 1326 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to